OpenStudy (anonymous):

The numThe number of cars an American family owns follows the distribution below: Number of Cars 0 1 2 3 4 5 Probability 0.09 0.36 0.35 0.13 0.05 0.02 a) Verify that this is a legitimate probability distribution. b) Interpret (in words) the notation P(X > 2). c) Interpret (in words) the notation P(X = 2). d) Find P(X > 2). (5 points)

ganeshie8 (ganeshie8):

for a distribution to be legitimate, it must satisfy below : 1) probabilities must add up to 1 2) each probability must be between [0, 1]

ganeshie8 (ganeshie8):

test whether the given distribution satisfies both conditions

OpenStudy (anonymous):

0.09+0.36+0.35+0.13+0.05+0.02=1

ganeshie8 (ganeshie8):

also each probability is between [0, 1] so its a legitimate distribution

ganeshie8 (ganeshie8):

for part b)

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

can u tell me what "X" represents here ?

ganeshie8 (ganeshie8):

"X" represents "number of cars an American family owns " right ?

ganeshie8 (ganeshie8):

So, wat does P(X > 2) represent ?

OpenStudy (anonymous):

not sure

ganeshie8 (ganeshie8):

give ur best try

ganeshie8 (ganeshie8):

think a bit and may be say something

OpenStudy (anonymous):

b. 0.09+0.36+0.35+0.13=0.93

OpenStudy (anonymous):

or 0.09+0.02=0.11

ganeshie8 (ganeshie8):

nope, we dont need to do any calculaitons for parts b and c

ganeshie8 (ganeshie8):

we just need to interpret what the given expression means, thats all

ganeshie8 (ganeshie8):

b) Interpret (in words) the notation P(X > 2).

ganeshie8 (ganeshie8):

wat does the P(X > 2) mean ?

OpenStudy (anonymous):

x is greater than 2

ganeshie8 (ganeshie8):

yes, wat does "X" represent here ?

OpenStudy (anonymous):

the numbers of cars

ganeshie8 (ganeshie8):

"X" represents " number of cars an American family owns "

ganeshie8 (ganeshie8):

so, P(X > 2) represents the "Probability for number of cars an American family owns is MORE THAN 2"

ganeshie8 (ganeshie8):

thats what P(X > 2) means okay ?

ganeshie8 (ganeshie8):

can u do part c similarly ?

ganeshie8 (ganeshie8):

c) Interpret (in words) the notation P(X = 2).

ganeshie8 (ganeshie8):

wat does P(X = 2) represent ?

OpenStudy (anonymous):

c. so P(X=2) the probability number of cars an american family owns equal to 2

ganeshie8 (ganeshie8):

you got it !

ganeshie8 (ganeshie8):

wat about part d ?

ganeshie8 (ganeshie8):

for this part, u need to do calculation

OpenStudy (anonymous):

d. 0.09+0.02=0.11

ganeshie8 (ganeshie8):

how ?

ganeshie8 (ganeshie8):

d) Find P(X > 2). (5 points)

ganeshie8 (ganeshie8):

P(X > 2) = P(X = 3) + P(X = 4) + P(C = 5) = ?

OpenStudy (anonymous):

add the first and last number or add 0.09 0.36+0.35+0.13=0.93

ganeshie8 (ganeshie8):

nope, btw why wud u do that ?

OpenStudy (anonymous):

ok

ganeshie8 (ganeshie8):

the question is asking us to find probability for having MORE THAN TWO cars

ganeshie8 (ganeshie8):

that means u need to add probabilities for having 3 or 4 or 5 cars ok ?

OpenStudy (anonymous):

0.35+0.13+0.05+0.02=.55

ganeshie8 (ganeshie8):

not quite, u should not add P(X = 2)

ganeshie8 (ganeshie8):

u need add oly 3 or 4 or 5 cars probabilities

OpenStudy (anonymous):

0.13+0.05+0.02=0.37

ganeshie8 (ganeshie8):

0.13 + 0.05 + 0.02 = 0.2

ganeshie8 (ganeshie8):

you should get 0.2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks

ganeshie8 (ganeshie8):

np :)